1. A sample of methane gas at a pressure of 0.809 atm and a temperature of 237 °C, occupies a volume of 417 mL. If the gas is heated at constant pressure until its volume is 666 mL, the temperature of the gas sample will be ___ °C.

2. A sample of krypton gas at a pressure of 958 mm Hg and a temperature of 65 °C, occupies a volume of 1683.9 mL. If the gas is cooled at constant pressure to a temperature of 26 °C, the volume of the gas sample, in liters, will be ___L.
1
Expert's answer
2019-10-22T05:43:48-0400

P*V=n*R*T

1) n, R, P =const : V1/V2=T1/T2

T1=237 + 273.15=510.15K

417/666 = 510.15/T2

T2 = 814. 77 K = 814.77-273.15 C = 541.62 C

2)n, R, P = const : V1/V2=T1/T2

T1=338.15K T2=299.15K

V2=1489.7 ml = 1.49L